# Downstream Task Guided Masking Learning in Masked Autoencoders Using Multi-Level Optimization

### v1

@@ Abstract
Masked Autoencoder (MAE) is a notable method for self-supervised pretraining in visual representation learning. It operates by randomly masking image patches and reconstructing these masked patches using the unmasked ones. A key limitation of MAE lies in its disregard for the varying informativeness of different patches, as it uniformly selects patches to mask. To overcome this, some approaches propose masking based on patch informativeness. However, these methods often do not consider the specific requirements of downstream tasks, potentially leading to suboptimal representations for these tasks. In response, we introduce the Multi-level Optimized Mask Autoencoder (MLO-MAE), a novel framework that leverages end-to-end feedback from downstream tasks to learn an optimal masking strategy during pretraining. Our experimental findings highlight MLO-MAE’s significant advancements in visual representation learning. Compared to existing methods, it demonstrates remarkable improvements across diverse datasets and tasks, showcasing its adaptability and efficiency. Our code is available at: https://github.com/Alexiland/MLOMAE
Machine Learning, ICML
@@ 1 Introduction
In the rapidly evolving field of self-supervised learning (Balestriero et al., 2023; Gui et al., 2023), particularly in visual representation learning, Masked Autoencoder (MAE) (He et al., 2022) has emerged as a prominent approach, which draws inspiration from the successful masked language models like BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019). Similar to how BERT learns textual representations by predicting randomly masked tokens, MAE is designed to learn visual representations by masking random patches of an image and then reconstructing them using the remaining unmasked patches.
@@
Although MAE has shown empirical success, it applies a uniform random approach to masking patches, overlooking the variable distribution of information across different image regions (Chen et al., 2023; Kong & Zhang, 2023; Wang et al., 2023; Liu et al., 2023). It assumes equal informativeness across all parts of an image, an assumption that does not always hold true. Some image areas (e.g., foreground objects) may hold more critical information than others (e.g., background), a factor not considered in MAE’s current design. Such oversight might hinder the model’s capability and efficiency in learning representations. This lack of distinction between more and less informative regions in the MAE could lead to disproportionate allocations of computational resources. Consequently, the model might spend excessive effort on less significant areas while inadequately processing and capturing the nuances in regions that contain more valuable information.
@@
To mitigate this limitation, various strategies have been suggested for masking patches contingent on their informativeness. Key approaches include masking regions with high attention scores to prioritize areas of interest (Li et al., 2021; Kakogeorgiou et al., 2022); employing semantic segmentation to identify and mask regions rich in information (Li et al., 2022; Chen et al., 2023); and utilizing adversarial models that generate masks to enhance learning efficiency (Shi et al., 2022). These approaches aim to refine the masking process by prioritizing patches based on the level of information they contain, rather than treating all patches uniformly.
@@
Although these methods are promising, they mask patches without incorporating feedback from downstream tasks. Their process involves two separate stages: initially employing a specific masking strategy to pretrain a model, then fine-tuning this model for downstream tasks, with the hope that the model pretrained using this strategy will be effective for these tasks. During this process, the design of the masking strategy is not influenced by the requirements of the downstream tasks. As a result, the representations developed through this strategy may not be well-aligned with the needs of these tasks, which could limit their effectiveness.
@@
To bridge this gap, we propose a downstream task guided masking strategy learning framework based on multi-level optimization (MLO) (Vicente & Calamai, 1994). Our approach utilizes feedback from downstream tasks to autonomously learn the optimal masking strategy. It integrates model pretraining with fine-tuning in an end-to-end manner, allowing the fine-tuning performance on downstream tasks to directly influence the masking process during pretraining. Our method learns a masking network to mask patches. It processes an input image to identify specific patches for masking. Our method consists of three interconnected stages. In the first stage, a preliminary version of the masking network masks certain patches, followed by the pretraining of an image encoder tasked with reconstructing these masked patches. In the second stage, we utilize this encoder to construct a downstream model, which is subsequently trained using the training dataset specific to a downstream task. The final stage involves evaluating the downstream model using a held-out validation dataset. The effectiveness of the masking network is indirectly measured by the downstream model’s validation performance. An inferior masking network might fail to correctly identify the optimal patches for masking, leading to ineffective pretraining of the image encoder. When applied to the downstream task, the encoder’s inadequate representation learning capabilities will lead to suboptimal validation performance of the downstream model. To prevent this, we continuously refine the masking network, ensuring it maximizes downstream validation performance. Each of these stages is formulated as one level of optimization problem in our MLO framework. The three levels of optimization problems are mutually dependent on each other and solved jointly. This enables the three stages to be conducted end-to-end, where the downstream validation performance closely guides the learning of the masking network.
@@
  • Our method showcases remarkable transfer learning abilities, in fine-grained classification and semantic segmentation tasks, demonstrated on datasets including CUB-200-2011, Stanford Cars, iNaturalist 2019, and ADE20K.
@@ 2.1 Masked Autoencoders
Following the success of masked language models in the field of natural language processing (Devlin et al., 2018), various masked image models have been proposed (Chen et al., 2020; Bao et al., 2022). Among them, Masked Autoencoder (MAE) has become a promising methodology for generic visual pretraining (He et al., 2022). MAE is a denoising autoencoder that randomly masks the input image and tries to reconstruct the missing pixels. It uses a high masking ratio (75% in MAE compared to 15% in BERT) and a lightweight decoder architecture that forces the encoder to learn meaningful visual representations. Zhang et al. (2022) propose a theoretical framework to understand the role of masking in MAE, and introduce a Uniformity-enhanced MAE (U-MAE) to address the dimensional collapse issue. Despite MAE’s effectiveness, recent works underscore the importance of replacing the random patch masking method in MAE with more sophisticated masking strategies (Kakogeorgiou et al., 2022; Shi et al., 2022). For instance, MST (Li et al., 2021) utilizes attention maps to guide the masking process, selectively obscuring less attended regions to maintain foreground information. SemMAE (Li et al., 2022) combines a StyleGAN-based decoder with the MAE decoder and leverages attention maps from the StyleGAN decoder to provide semantic cues for patch masking. Furthermore, some recent methods propose to use a learnable masking module to generate masking strategies and optimize the masking module in the pretraining process. For example, AutoMAE (Chen et al., 2023) links a differentiable mask generator with MAE using Gumbel-Softmax (Jang et al., 2016), following a similar two-stage setup as in SemMAE. CL-MAE (Madan et al., 2024) leverages curriculum learning to enhance MAE by progressively increasing the complexity of the masks generated from a learnable masking module. Compared to these existing methods, the key distinction of our approach lies in the utilization of feedback from downstream tasks to inform the development of masking strategies, a mechanism absent in the current methodologies.
Figure 1: An overview of MLO-MAE, which consists of three stages performed end-to-end. Modules with learnable parameters are indicated in blue, and those with frozen parameters are in orange.
@@ 3.1 Overview
We introduce the Multi-level Optimized MAE (MLO-MAE), a self-supervised visual representation learning method that automatically learns an optimal masking strategy in a Masked Autoencoder. As illustrated in Figure 1, the MLO-MAE architecture comprises three key components: a masking network $T$, a Vision Transformer (ViT) (Dosovitskiy et al., 2020) based image encoder $E$, and a classification head $C$. Given an input image, the masking network identifies the patches to be masked. The image encoder extracts a representation for an input image. It is pretrained by reconstructing masked patches from unmasked ones. The classification head $C$ predicts a class label from an image representation extracted by the encoder. MLO-MAE leverages an image classification dataset $\mathcal{D}$, comprising pairs of images and their corresponding class labels. This dataset is divided into a training subset $\mathcal{D}^{tr}$ and a validation subset $\mathcal{D}^{val}$. The image encoder is pretrained on unlabeled input images (denoted as $\mathcal{D}_{u}$) from $\mathcal{D}^{tr}$ (excluding labels).
Our method operates in three end-to-end stages. In the first stage, a tentative version of the masking network masks input images from $\mathcal{D}_{u}$. Then the image encoder is pretrained on these masked images. In the second stage, the pretrained encoder extracts representations for images in the training set $\mathcal{D}^{tr}$. These representations, along with their corresponding labels, are used to train the classification head by minimizing classification losses. In the third stage, the pretrained encoder extracts representations for images in the validation set $\mathcal{D}^{val}$. The trained classification head then uses these representations to predict labels. Validation loss is computed by comparing these predictions with the actual labels in $\mathcal{D}^{val}$. As shown later, the validation loss is a function of the masking network $T$. This loss serves as a metric for assessing the effectiveness of $T$. To enhance the functionality of $T$, we focus on minimizing this loss.
Our method employs a multi-level optimization approach, involving several nested optimization problems. Optimal parameters obtained at each lower level serve as inputs for the loss functions at the subsequent upper levels. Conversely, non-optimal parameters from the upper levels are utilized to define the loss functions at lower levels. Each of the three aforementioned stages corresponds to a single level of optimization. Different stages are executed end-to-end by solving problems at different levels jointly.
@@ 3.2 Multi-level Optimization Framework
Given an input image $X\in D_{u}$ divided into $N$ non-overlapping patches of equal size, denoted as $\{P_{i}\}_{i=1}^{N}$, the masking network $T$ takes $X$ as input and generates a probability $\sigma(P_{i},X;T)$ for each patch $P_{i}$, which represents the likelihood that $P_{i}$ should be masked. Given a masking ratio $r$, a hyperparameter dictating the proportion of patches to be masked, we first rank all patches in descending order based on their masking probabilities. We then select the top $N\times r$ patches denoted as $\mathcal{M}(X;T,r)$, those with the highest probabilities, and mask them. The remaining patches, denoted as $X-\mathcal{M}(X;T,r)$, are unmasked.
Then we feed the unmasked patches into an autoencoder (He et al., 2022), which consists of the image encoder $E$ and a decoder $D$, to reconstruct the masked patches $\mathcal{M}(X;T,r)$. In detail, the unmasked patches are first processed by the image encoder $E$, which is responsible for extracting their representations. Then, these representations are input into the decoder $D$. The decoder’s role is to accurately predict the pixel values of the masked patches. To evaluate the performance of this reconstruction, we employ a reconstruction loss, $\mathcal{L}_{rec}$, defined as the squared differences between the predicted and ground truth pixel values of the masked patches. Importantly, the reconstruction loss for the $j$-th masked patch $\mathcal{M}_{j}(X;T,r)$ is weighted according to its masking probability $\sigma(\mathcal{M}_{j}(X;T,r),X;T)$. This probability reflects the likelihood of a patch being masked and thus, guides the autoencoder to prioritize the reconstruction of patches deemed more likely to be masked.
@@
$\small\begin{array}[]{l}E^{*}(T),D^{*}=\underset{E,D}{\text{argmin}}\sum\limits_{X\in\mathcal{D}_{u}}\sum\limits_{j=1}^{N\times r}\sigma(\mathcal{M}_{j}(X;T,r),X;T)\times\\ \qquad\qquad\quad\mathcal{L}_{rec}(X-\mathcal{M}(X;T,r),\mathcal{M}_{j}(X;T,r);E,D).\end{array}$ (1)
@@
Inspired by (Liu et al., 2018), we develop an efficient hypergradient-based method to solve the problem in Eq.(4). First, we approximate the optimal solutions $E^{*}(T)$ and $D^{*}$ by executing several iterations (termed as unrolling steps) of gradient descent updates of $E$ and $D$ against the loss function at the first level. The approximation of $E^{*}(T)$ is then plugged into the second-level loss, and $C^{*}(E^{*}(T))$ is similarly approximated using multiple steps of gradient descent updates of $C$ against this approximate loss. The approximations of $E^{*}(T)$ and $C^{*}(E^{*}(T))$ are then applied to the third-level loss, enabling the gradient descent update of $T$. This iterative process of updating continues until convergence is achieved.
Table 1: Top-1 accuracy (%) on the test sets of CIFAR-10, CIFAR-100, and ImageNet, in fine-tuning experiments. Results marked with * are taken from U-MAE (Zhang et al., 2022) and AutoMAE (Chen et al., 2023). The baseline methods SemMAE and AutoMAE are not included in the comparison on CIFAR-10 and CIFAR-100, due to the absence of reported results for these datasets in their original publications and the unavailability of their implementation code for conducting evaluations on these datasets.
@@
ImageNet-1K 77.9 $**{}^{\text{*}}$82.9 $**{}^{\text{*}}$83.0 $**{}^{\text{*}}$83.3 $**{}^{\text{*}}$83.3 84.8
4 Experiments
4.1 Datasets
We evaluated our MLO-MAE method on three benchmark image classification datasets: CIFAR-10 (Krizhevsky et al., a), CIFAR-100 (Krizhevsky et al., b), and ImageNet-1K (Deng et al., 2009). CIFAR-10 and CIFAR-100 contain 50K training images and 10K test images from 10 and 100 classes, respectively. ImageNet contains 1.3M training images and 50K validation images (used as test set) from 1000 classes. In our approach, we divide the original training set of each dataset into two subsets with a ratio of 8:2, which are used as the $D^{tr}$ and $D^{val}$ in Eq.(4), respectively. Unlabeled images $D_{u}$ used in MLO-MAE’s first stage are from $D^{tr}$. Baseline methods are trained on the unsplit original training set. It is important to note that our method does not unfairly utilize more data than the baselines. To assess the model’s transferability to fine-grained classification tasks, we further tested it on the CUB-200-2011 dataset (Wah et al., 2011), Stanford Cars dataset (Krause et al., 2013), and iNaturalist 2019 dataset (Van Horn et al., 2018). Additionally, we evaluated MLO-MAE’s semantic segmentation capabilities using the ADE20K dataset (Zhou et al., 2017). For comprehensive details about these datasets, please refer to Appendix B.1. In all experiments, performance on the test set is reported.
4.2 Experimental Settings
Model Setup and Hyperparameters.
@@ 4.2 Experimental Settings
We conducted comparisons with several baselines, including: 1) vanilla MAE (He et al., 2022) and U-MAE (Zhang et al., 2022), which employ uniform random masking of images; 2) SemMAE (Li et al., 2022) and AutoMAE (Chen et al., 2023), which mask patches according to their informativeness; and 3) the Vision Transformer (ViT) (Dosovitskiy et al., 2020) without pretraining by MAE methods (i.e., directly trained for classification from scratch). ViT-B was used as the image encoder in these methods.

### v2

@@ Abstract
Masked Autoencoder (MAE) is a notable method for self-supervised pretraining in visual representation learning. It operates by randomly masking image patches and reconstructing these masked patches using the unmasked ones. A key limitation of MAE lies in its disregard for the varying informativeness of different patches, as it uniformly selects patches to mask. To overcome this, some approaches propose masking based on patch informativeness. However, these methods often do not consider the specific requirements of downstream tasks, potentially leading to suboptimal representations for these tasks. In response, we introduce the Multi-level Optimized Mask Autoencoder (MLO-MAE), a novel framework that leverages end-to-end feedback from downstream tasks to learn an optimal masking strategy during pretraining. Our experimental findings highlight MLO-MAE’s significant advancements in visual representation learning. Compared to existing methods, it demonstrates remarkable improvements across diverse datasets and tasks, showcasing its adaptability and efficiency. Our code is available at https://github.com/Alexiland/MLO-MAE
[…]
@@ 1 Introduction
In the rapidly evolving field of self-supervised learning (Balestriero et al., 2023; Gui et al., 2023), particularly in visual representation learning, Masked Autoencoder (MAE) (He et al., 2022) has emerged as a prominent approach, which draws inspiration from the successful masked language models like BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019). Similar to how BERT learns textual representations by predicting randomly masked tokens, MAE is designed to learn visual representations by masking random patches of an image and then reconstructing them using the remaining unmasked ones.
@@
Although MAE has shown empirical success, it applies a uniform random approach to mask patches, overlooking the varying distribution of information across different image regions (Chen et al., 2023; Kong & Zhang, 2023; Wang et al., 2023; Liu et al., 2023). It assumes equal informativeness across all parts of an image, an assumption that does not always hold true. Some image areas may hold more critical information than others, a factor not considered in MAE’s current design. Such oversight might hinder the model’s capability and efficiency in learning representations. This lack of distinction between more and less informative regions in the MAE could lead to disproportionate allocations of computational resources. Consequently, the model might spend excessive effort on less significant areas while inadequately processing and capturing the nuances in regions that contain more valuable information.
@@
To mitigate this limitation, various strategies have been suggested for masking patches contingent on their informativeness. Key approaches include masking regions with high attention scores to prioritize areas of interest (Li et al., 2021; Kakogeorgiou et al., 2022); employing semantic segmentation to identify and mask regions rich in information (Li et al., 2022); automatically learning a masking module (Madan et al., 2024); and learning a differentiable mask generator via adversarial training (Chen et al., 2023). These approaches aim to refine the masking process by prioritizing patches based on the level of information they contain, rather than treating all patches uniformly.
@@
Although these methods are promising, they mask patches without incorporating feedback from downstream tasks. Their process involves two separate stages: initially employing a specific masking strategy to pretrain an image encoder, then using this encoder to perform downstream tasks (via finetuning (He et al., 2022) or linear probing (He et al., 2022)), with the hope that the encoder pretrained using this strategy will be effective for these tasks. During this process, the design of the masking strategy is not influenced by the requirements of the downstream tasks. As a result, the representations developed through this strategy may not be well-aligned with the needs of these tasks, which could limit their effectiveness.
@@
To bridge this gap, we propose a downstream task guided masking strategy learning framework based on multi-level optimization (MLO) (Vicente & Calamai, 1994). Our approach utilizes feedback from downstream tasks to autonomously learn the optimal masking strategy. It pretrains an image encoder and applies the pretrained encoder to perform a downstream task in an end-to-end manner, allowing the downstream task’s performance to directly influence the masking process during pretraining. Our method learns a masking network to mask patches. It processes an input image to identify specific patches for masking. Our method consists of three interconnected stages. In the first stage, a preliminary version of the masking network masks certain patches, followed by the pretraining of an image encoder tasked with reconstructing these masked patches. In the second stage, we utilize this encoder to construct a downstream model, which is subsequently trained using the training dataset specific to a downstream task. The final stage involves evaluating the downstream model using a held-out validation dataset. The effectiveness of the masking network is indirectly measured by the downstream model’s validation performance. An inferior masking network might fail to correctly identify the optimal patches for masking, leading to ineffective pretraining of the image encoder. When applied to the downstream task, the encoder’s inadequate representation learning capabilities will lead to suboptimal validation performance of the downstream model. To prevent this, we continuously refine the masking network, ensuring it maximizes downstream validation performance. Each of these stages is formulated as one level of optimization problem in our MLO framework. The three levels of optimization problems are mutually dependent on each other and solved jointly. This enables the three stages to be conducted end-to-end, where the downstream validation performance closely guides the learning of the masking network.
@@
  • Our method showcases remarkable transfer learning abilities, in fine-grained classification, semantic segmentation, and object detection tasks, demonstrated on datasets including CUB-200-2011, Stanford Cars, iNaturalist 2019, ADE20K, and MS-COCO.
@@ 2.1 Masked autoencoders
Following the success of masked language models in the field of natural language processing (Devlin et al., 2018), various masked image models have been proposed (Chen et al., 2020; Bao et al., 2022). Among them, Masked Autoencoder (MAE) has become a promising methodology for generic visual pretraining (He et al., 2022). MAE is a denoising autoencoder that randomly masks the input image and tries to reconstruct the missing pixels. It uses a high masking ratio (75% in MAE compared to 15% in BERT) and a lightweight decoder architecture that forces the encoder to learn meaningful visual representations. Zhang et al. (2022) propose a theoretical framework to understand the role of masking in MAE, and introduce a Uniformity-enhanced MAE (U-MAE) to address the dimensional collapse issue. Despite MAE’s effectiveness, recent works underscore the importance of replacing the random patch masking method in MAE with more sophisticated masking strategies (Kakogeorgiou et al., 2022; Shi et al., 2022). For instance, MST (Li et al., 2021) utilizes attention maps to guide the masking process, selectively obscuring less attended regions to maintain important information. SemMAE (Li et al., 2022) combines a StyleGAN-based decoder with the MAE decoder and leverages attention maps from the StyleGAN decoder to provide semantic cues for patch masking. Furthermore, some recent methods propose to use a learnable masking module to generate masking strategies and optimize the masking module in the pretraining process. For example, AutoMAE (Chen et al., 2023) links a differentiable mask generator with MAE using Gumbel-Softmax (Jang et al., 2016), following a similar two-stage setup as in SemMAE. CL-MAE (Madan et al., 2024) leverages curriculum learning to enhance MAE by progressively increasing the complexity of the masks generated from a learnable masking module. Compared to these existing methods, the key distinction of our approach lies in the utilization of feedback from downstream tasks to inform the development of masking strategies, a mechanism absent in the current methodologies.
Figure 1: An overview of MLO-MAE, which consists of three stages performed end-to-end. Modules with learnable parameters are indicated in orange, and those with frozen parameters are in blue.
@@ 3.1 Overview
We introduce the Multi-level Optimized MAE (MLO-MAE), an end-to-end visual representation learning method which leverages the guidance from a downstream task to learn an optimal masking strategy automatically. Conceptually, MLO-MAE reimagines the visual representation learning by coupling pretraining with downstream task performance. MLO-MAE introduces a learnable masking strategy that prioritizes patches based on their relevance to a downstream task. This enables the model to focus its representation learning on features that are most critical for achieving better downstream performance.
As illustrated in Figure 1, the architecture of MLO-MAE consists of three key components: a Vision Transformer (ViT) (Dosovitskiy et al., 2020)-based image encoder $E$, a classification head $C$, and a masking network $T$. The masking network processes input images to identify patches for masking. The image encoder then extracts a representation by reconstructing the masked patches using information from the unmasked ones, serving as the backbone for visual representation learning. The classification head $C$ predicts a class label based on the image representation extracted by the encoder. The pretraining process is carried out on an unlabeled dataset $\mathcal{D}_{u}$, similar to other self-supervised methods. However, the learning of the masking network is guided by a downstream image classification task using a labeled dataset $\mathcal{D}$, which is further divided into a training subset $\mathcal{D}^{tr}$ and a validation subset $\mathcal{D}^{val}$.
MLO-MAE operates in three interconnencted stages. In the Stage I, the masking network $T$ generates a preliminary mask for input images from $\mathcal{D}_{u}$, and the image encoder $E$ is pretrained on these masked images by minimizing reconstruction loss. In the Stage II, the pretrained encoder extracts representations for images in the training subset $\mathcal{D}^{tr}$. These representations, along with their associated labels, are used to train the classification head $C$ by minimizing classification loss. Finally, in Stage III, the pretrained encoder extracts representations for the validation subset $\mathcal{D}^{val}$, which are passed to the classification head for label predictions. Validation loss, computed by comparing the predicted labels with the actual labels, serves as a feedback mechanism for evaluating the effectiveness of the masking network $T$. By focusing on minimizing this validation loss, the masking network $T$ is iteratively refined.
To seamlessly integrate these stages, MLO-MAE employs a multi-level optimization (MLO) framework, where each stage corresponds to one level of optimization. Optimal parameters obtained at each lower level serve as inputs for the loss functions at the subsequent upper levels. Conversely, non-optimal parameters from the upper levels are utilized to define the loss functions at lower levels. These nested optimization problems are solved jointly, enabling the three stages to operate in an end-to-end manner. This multi-level approach ensures that the pretraining process dynamically aligns with the downstream task requirements, allowing MLO-MAE to learn a masking strategy that optimally enhances visual representation learning.
@@ 3.2 Multi-level optimization framework
Given an input image $X\in D_{u}$ divided into $N$ non-overlapping patches of equal size, denoted as $\{P_{i}\}_{i=1}^{N}$, the masking network $T$ takes $X$ as input and generates a probability $\sigma(P_{i},X;T)$ for each patch $P_{i}$, which represents the likelihood that $P_{i}$ should be masked. Given a masking ratio $r$, a hyperparameter dictating the proportion of patches to be masked, we first rank all patches in descending order based on their masking probabilities. We then select the top $N\times r$ patches denoted as $\mathcal{M}(X;T,r)$, those with the highest probabilities, and mask them. The remaining patches, denoted as $X-\mathcal{M}(X;T,r)$, are unmasked. Then we feed the unmasked patches into an autoencoder (He et al., 2022), which consists of the image encoder $E$ and a decoder $D$, to reconstruct the masked patches $\mathcal{M}(X;T,r)$. In detail, the unmasked patches are first processed by the image encoder $E$, which is responsible for extracting their representations. Then, these representations are input into the decoder $D$. The decoder’s role is to accurately predict the pixel values of the masked patches. To evaluate the performance of this reconstruction, we employ a reconstruction loss, $\mathcal{L}_{rec}$, defined as the squared differences between the predicted and ground truth pixel values of the masked patches. Importantly, the reconstruction loss for the $j$-th masked patch $\mathcal{M}_{j}(X;T,r)$ is weighted according to its masking probability $\sigma(\mathcal{M}_{j}(X;T,r),X;T)$. This probability reflects the likelihood of a patch being masked and thus, guides the autoencoder to prioritize the reconstruction of patches deemed more likely to be masked.
@@
$\small\begin{array}[]{l}E^{*}(T),D^{*}=\underset{E,D}{\text{argmin}}\sum\limits_{X\in\mathcal{D}_{u}}\sum\limits_{j=1}^{N\times r}\sigma(\mathcal{M}_{j}(X;T,r),X;T)\mathcal{L}_{rec}(X-\mathcal{M}(X;T,r),\mathcal{M}_{j}(X;T,r);E,D).\end{array}$ (1)
@@
Inspired by Liu et al. (2018), we develop an efficient hypergradient-based method to solve the problem in Eq.(4). First, we approximate the optimal solutions $E^{*}(T)$ and $D^{*}$ by executing several iterations (termed as unrolling steps) of gradient descent updates of $E$ and $D$ against the loss function at the first level. The approximation of $E^{*}(T)$ is then plugged into the second-level loss, and $C^{*}(E^{*}(T))$ is similarly approximated using multiple steps of gradient descent updates of $C$ against this approximate loss. The approximations of $E^{*}(T)$ and $C^{*}(E^{*}(T))$ are then applied to the third-level loss, enabling the gradient descent update of $T$. This iterative process of updating continues until convergence is achieved. Details of this optimization algorithm are deferred to Appendix A and Algorithm 1.
4 Experiments
4.1 Experimental settings
Model setup and hyperparameters.
Table 1: Top-1 accuracy (%) on the test sets of CIFAR-10, CIFAR-100, and ImageNet, in fine-tuning experiments. The baseline methods SemMAE and AutoMAE are not included in the comparison on CIFAR-10 and CIFAR-100, due to the absence of reported results for these datasets in their original publications and the unavailability of their implementation code for conducting evaluations on these datasets.
@@ 4.1 Experimental settings
ImageNet-1K 77.9 83.6 83.0 83.3 83.3 84.8
@@
We conducted comparisons with several baselines, including: 1) vanilla MAE (He et al., 2022) and U-MAE (Zhang et al., 2022), which employ uniform random masking of images; 2) SemMAE (Li et al., 2022) and AutoMAE (Chen et al., 2023), which mask patches according to their informativeness; and 3) the Vision Transformer (ViT) (Dosovitskiy et al., 2020) without pretraining by MAE methods (i.e., directly trained for classification from scratch). ViT-B was used as the image encoder in these methods. Following their original papers, the number of pre-training epochs for MAE, U-MAE, SemMAE, and AutoMAE on ImageNet are 1600, 200, 800, and 800 respectively. The patch size in all methods is 16 for ImageNet.
@@
It is important to note that our method does not unfairly utilize more labeled data than the baselines. The labeled data used in Stage II and III of our framework is identical to that used in the fine-tuning phrase of the baselines.